# Neural Incremental Input-to-State Stable Control Lyapunov Functions for Unknown Continuous-time Systems

### v1

@@ 5.3 Jet Engine
We consider a nonlinear Moore-Grietzer Jet Engine Model in no-stall mode [19], whose dynamics is governed by the following set of equations.
@@
where $\mathsf{x}_{1}=\mu-1,\mathsf{x}_{2}=\zeta-\rho-2$ with $\mu,\zeta,\rho$ denote the mass flow, the pressure rise and a constant respectively. $\tau=0.01$ is the sampling time. We consider the state space of the system to be ${\mathbf{X}}=[-0.25,0.25]\times[-0.25,0.25]$. Moreover, we consider the input set to be bounded within ${\mathbf{W}}=[-0.25,0.25]$. Also, we consider the model to be unknown. However, we estimate the Lipschitz constants $\mathcal{L}_{x}=1.213,\mathcal{L}_{u}=1$.
@@ 5.4 Rotating Spacecraft Model
We consider another example of a rotating rigid spacecraft model [18], whose discrete-time dynamics is governed by the following set of equations.
@@
where $\mathsf{x}=[\mathsf{x}_{1},\mathsf{x}_{2},\mathsf{x}_{3}]^{\top}$ denotes angular velocities $\omega_{1},\omega_{2},\omega_{3}$ along the principal axes respectively, $\mathsf{u}=[\mathsf{u}_{1},\mathsf{u}_{2},\mathsf{u}_{3}]^{\top}$ represents the torque input, and with $J_{1}=200,J_{2}=200,J_{3}=100$ denote the principal moments of inertia. $\tau=0.01$ is the sampling time. We consider the state space of the system to be ${\mathbf{X}}=[-0.25,0.25]\times[-0.25,0.25]\times[-0.25,0.25]$. Moreover, we consider the input set to be bounded within ${\mathbf{W}}=[-10,10]$. Note that the internal input has 3 dimensions, while the external input has a single dimension, that is, ${\mathbf{U}}\subset{\mathbb{R}}^{3}$, but ${\mathbf{W}}\subset{\mathbb{R}}$. In addition, we consider the model to be unknown. However, we estimate the Lipschitz constants $\mathcal{L}_{x}=0.160,\mathcal{L}_{u}=0.015$.
@@
One can see from Figure 4(a), 4(b), and 4(c), under the different input conditions, the trajectories corresponding to various states starting from different initial conditions maintain the same distance after some time instances under the influence of the controller.
@@ 5.5 Discussion and Comparisons
The proposed approach to synthesizing controllers, ensuring incremental stability for closed-loop continuous-time systems, provides several benefits. The data-driven neural controller synthesis utilizes the data from the state-space of the system. Existing approaches for designing controllers that ensure incremental input-to-state stability, such as those in [43, 15], require full knowledge of the system dynamics and assume a specific structure, typically a strict-feedback form. A recent work [33] proposed estimating the strict feedback nature using the Gaussian process and designing a backstepping-like control to make the system incrementally stable. Compared to these methods, our approach can handle any general nonaffine nonlinear system, as seen in the Case Study 5.1 of a general nonlinear system. Although recent advancement [38] uses data collected from system trajectories to design a controller for an input-affine nonlinear system with polynomial dynamics, our method extends to arbitrary nonlinear dynamics, as seen in the first case study, where neither input-affine structures nor polynomial forms are present.
The final case study demonstrates that the system possesses three control signals, but can be effectively controlled using a reduced number of external inputs. Thereby, the proposed methodology is able to make the closed-loop system controllable with a reduced number of inputs, simplifying the control implementation. This will be particularly beneficial in improving the scalability of abstraction-based controller synthesis approaches [30, 28, 16]. As discussed in [42], the computational complexity of constructing an input-based abstraction of an incrementally stable system depends on the input dimension of the system. As we are able to make the system incrementally stable with a reduced number of input dimensions, the complexity of the abstraction is significantly reduced. This facilitates the efficient synthesis of additional formally verified controllers to meet other specifications for the closed-loop system.
@@ 6 Conclusion and Future Work
The work can be extended into constructing $\delta$ISS Control Lyapunov functions under the action of the controller for interconnected systems as well, which is currently being investigated. We will try to synthesize $\delta$-ISS Lyapunov functions for each subsystem and claim how they can be used to formally guarantee the incremental stability of the large system. In addition, future direction of this work can be in the purview of stochastic systems, providing a formal guarantee of their incremental stability.
@@
The proof is inspired by the proof of Theorem 2.6 of [40]. Let $\mathsf{x}_{x,\mathsf{w}}(t)$ denotes the state of the closed-loop system at time $t$ under input signal $\mathsf{w}$ from the initial condition $x=\mathsf{x}_{x,\mathsf{w}}(0)$. Note that, since the system is forward invariant under the action of the controller $g_{\bar{\theta},\bar{b}}$, for all $t\geq 0$ we have $\mathsf{x}_{x,\mathsf{w}}(t)\in{\mathbf{X}}$. Consider the closed-loop system admits a $\delta$-ISS-CLF under a controller. Now, by using property (i) of Definition 2.5, we get
@@
for any $t\in{\mathbb{R}}_{0}^{+}$. Now using condition (ii) of Definition 2.5 and the comparison lemma [18, Lemma 3.4], we get
@@
$\leq\alpha_{1}^{-1}\bigg{(}e^{-\kappa t}V(x,\hat{x})+\frac{1-e^{-\kappa t}}{\kappa}\sigma(\lVert\mathsf{w}-\hat{\mathsf{w}}\rVert)\bigg{)}$
$\leq\alpha_{1}^{-1}\bigg{(}e^{-\kappa t}V(x,\hat{x})+\frac{1}{\kappa}\sigma(\lVert\mathsf{w}-\hat{\mathsf{w}}\rVert)\bigg{)}$
@@
where $\Psi(\rho,\Upsilon):=\alpha_{1}^{-1}(\rho+\Upsilon),\rho:=e^{-\kappa t}\alpha_{2}(|x-\hat{x}|),\Upsilon:=\frac{1}{\kappa}\sigma(\lVert\mathsf{w}-\hat{\mathsf{w}}\rVert)$. Notice that $\Psi$ is nondecreasing in both variables, we get:
$|\mathsf{x}_{x,\mathsf{w}}(t)-\mathsf{x}_{\hat{x},\hat{\mathsf{w}}}(t)|\leq\xi(e^{-\kappa t}\alpha_{2}(|x-\hat{x}|))+\xi(\frac{1}{\kappa}\sigma(\lVert\mathsf{w}-\hat{\mathsf{w}}\rVert)),$
@@
$|\mathsf{x}_{x,\mathsf{w}}(t)-\mathsf{x}_{\hat{x},\hat{\mathsf{w}}}(t)|$ $\leq\alpha_{1}^{-1}(2e^{-\kappa t}\alpha_{2}(|x-\hat{x}|))+\alpha_{1}^{-1}(\frac{2}{\kappa}\sigma(\lVert\mathsf{w}-\hat{\mathsf{w}}\rVert)).$
@@
$\beta(|x-\hat{x}|,t):=\alpha_{1}^{-1}(2e^{-\kappa t}\alpha_{2}(|x-\hat{x}|)),$ (23)
$\gamma(\lVert\mathsf{w}-\hat{\mathsf{w}}\rVert):=\alpha_{1}^{-1}(\frac{2}{\kappa}\sigma(\lVert\mathsf{w}-\hat{\mathsf{w}}\rVert)),$ (24)
@@
Appendix B Proof of Theorem 3.2
We will use the following Lemma to prove the Theorem.
Lemma B.1
[18, Exercise 3.3] If two functions $f_{1}$ and $f_{2}$ are Lipschitz continuous with constants $\mathbf{L}_{1}$ and $\mathbf{L}_{2}$ respectively and they are bounded, i.e., $\sup|f_{1}|\leq\mathcal{M}_{1}$ and $\sup|f_{2}|\leq\mathcal{M}_{2}$, then their product $f_{1}f_{2}$ is also Lipschitz continuous with Lipschitz constant $\mathcal{M}_{1}\mathbf{L}_{2}+\mathcal{M}_{2}\mathbf{L}_{1}$.
Here, we demonstrate under condition (10), the obtained $\delta$-ISS-CLF and the controller from SCP satisfy the conditions of Definition 2.5. The optimal $\eta_{S}^{*}$, obtained through solving the (9), guarantees for any $x_{q},x_{r}\in\mathcal{X},w_{q},w_{r}\in\mathcal{W}$, we have
$-V_{\theta,b}(x_{q},x_{r})+k_{1}|x_{q}-x_{r}|^{\gamma_{1}}\leq\eta_{S}^{*},$
$V_{\theta,b}(x_{q},x_{r})-k_{2}|x_{q}-x_{r}|^{\gamma_{2}}\leq\eta_{S}^{*},$
$\frac{\partial V_{\theta,b}}{\partial x_{q}}f(x_{q},g_{\bar{\theta},\bar{b}}(x_{q},w_{q}))+\frac{\partial V_{\theta,b}}{\partial x_{r}}f(x_{r},g_{\bar{\theta},\bar{b}}(x_{r},w_{r}))+\kappa V_{\theta,b}(x_{q},x_{r})-k_{w}|w_{q}-w_{r}|^{\gamma_{w}}\leq\eta_{S}^{*},$
$-\frac{\partial h}{\partial x}\Bigr{\rvert}_{x=x_{q}}(f(x_{q},g_{\bar{\theta},\bar{b}}(x_{q},w_{q})))-\mu_{h}h(x_{q})\leq\eta_{S}^{*}.$
Now, from (7), we infer that $\forall x\in{\mathbf{X}},\text{there exists}$ $x_{r}$ s.t. $|x-x_{r}|\leq\varepsilon$ as well as $\forall w\in{\mathcal{W}},\text{there exists}$ $w_{r}$ s.t. $|w-w_{r}|\leq\varepsilon$. Hence, $\forall x,\hat{x}\in{\mathbf{X}},\forall w,\hat{w}\in{\mathcal{W}}$:
(a) $-V_{\theta,b}(x,\hat{x})+k_{1}|x-\hat{x}|^{\gamma_{1}}$
$=\big{(}-V_{\theta,b}(x,\hat{x})+V_{\theta,b}(x_{q},x_{r})\big{)}+\big{(}-V_{\theta,b}(x_{q},x_{r})+k_{1}|x_{q}-x_{r}|^{\gamma_{1}}\big{)}+\big{(}-k_{1}|x_{q}-x_{r}|^{\gamma_{1}}+k_{1}|x-\hat{x}|^{\gamma_{1}}\big{)}$
$\leq\mathcal{L}_{L}|(x,\hat{x})-(x_{q},x_{r})|+\eta_{S}^{*}+2\mathsf{L}_{1}\varepsilon$
$\leq\sqrt{2}\mathcal{L}_{L}\varepsilon+\eta_{S}^{*}+2\mathsf{L}_{1}\varepsilon\leq\mathcal{L}\varepsilon+\eta_{S}^{*}\leq 0.$
(b) $V_{\theta,b}(x,\hat{x})-k_{2}|x-\hat{x}|^{\gamma_{2}}$
$=\big{(}V_{\theta,b}(x,\hat{x})-V_{\theta,b}(x_{q},x_{r})\big{)}+\big{(}V_{\theta,b}(x_{q},x_{r})-k_{2}|x_{q}-x_{r}|^{\gamma_{2}}\big{)}+\big{(}k_{2}|x_{q}-x_{r}|^{\gamma_{2}}-k_{2}|x-\hat{x}|^{\gamma_{2}}\big{)}$
$\leq\mathcal{L}_{L}|(x,\hat{x})-(x_{q},x_{r})|+\eta_{S}^{*}+2\mathsf{L}_{2}\varepsilon$
$\leq\sqrt{2}\mathcal{L}_{L}\varepsilon+\eta_{S}^{*}+2\mathsf{L}_{2}\varepsilon\leq\mathcal{L}\varepsilon+\eta_{S}^{*}\leq 0.$
(c) $\frac{\partial V_{\theta,b}}{\partial x}f(x,g_{\bar{\theta},\bar{b}}(x,w))+\frac{\partial V_{\theta,b}}{\partial\hat{x}}f(\hat{x},g_{\bar{\theta},\bar{b}}(\hat{x},\hat{w}))+\kappa V_{\theta,b}(x,\hat{x})-k_{w}|w-\hat{w}|^{\gamma_{w}}$
=$=\{\frac{\partial V_{\theta,b}}{\partial x}f(x,g_{\bar{\theta},\bar{b}}(x,w)-\frac{\partial V_{\theta,b}}{\partial x_{q}}f(x_{q},g_{\bar{\theta},\bar{b}}(x_{q},w_{q}))\}+\frac{\partial V_{\theta,b}}{\partial x_{q}}f(x_{q},g_{\bar{\theta},\bar{b}}(x_{q},w_{q}))+\{\frac{\partial V_{\theta,b}}{\partial\hat{x}}f(\hat{x},g_{\bar{\theta},\bar{b}}(\hat{x},\hat{w}))$
$\quad-\frac{\partial V_{\theta,b}}{\partial x_{r}}f(x_{r},g_{\bar{\theta},\bar{b}}(x_{r},w_{r}))\}+\frac{\partial V_{\theta,b}}{\partial x_{r}}f(x_{r},g_{\bar{\theta},\bar{b}}(x_{r},w_{r}))+\kappa\{V_{\theta,b}(x,\hat{x})-V_{\theta,b}(x_{q},x_{r})\}+\kappa V_{\theta,b}(x_{q},x_{r})$
$\quad-\{k_{w}|w-\hat{w}|^{\gamma_{w}}-k_{w}|w_{q}-w_{r}|^{\gamma_{w}}\}-k_{w}|w_{q}-w_{r}|^{\gamma_{w}}$
$\leq\eta_{S}^{*}+\sqrt{2}\kappa\mathcal{L}_{L}\varepsilon+2L_{u}\varepsilon+2\{\frac{\partial V_{\theta,b}}{\partial x}f(x,g_{\bar{\theta},\bar{b}}(x,w)-\frac{\partial V_{\theta,b}}{\partial x_{q}}f(x_{q},g_{\bar{\theta},\bar{b}}(x_{q},w_{q}))\}$
$\leq\eta_{S}^{*}+\sqrt{2}\kappa\mathcal{L}_{L}\varepsilon+2L_{u}\varepsilon+2(\mathcal{M}_{f}\mathcal{L}_{dL}+\mathcal{M}_{L}(\mathcal{L}_{x}+\sqrt{2}\mathcal{L}_{u}\mathcal{L}_{C}))\varepsilon$
$\leq\eta_{S}^{*}+(\sqrt{2}\kappa\mathcal{L}_{L}+2L_{u}+2(\mathcal{M}_{f}\mathcal{L}_{dL}+\mathcal{M}_{L}(\mathcal{L}_{x}+\sqrt{2}\mathcal{L}_{u}\mathcal{L}_{C})))\varepsilon\leq\eta_{S}^{*}+\mathcal{L}\varepsilon\leq 0.$
From Assumption 4, the gradient $\frac{\partial V_{\theta,b}}{\partial x}$ is bounded by $\mathcal{L}_{dL}$. Additionally, the function $f(x,g_{\bar{\theta},\bar{b}}(x,w))$ is Lipschitz continuous with a constant $\mathcal{L}_{x}+\sqrt{2}\mathcal{L}_{u}\mathcal{L}_{C}$. Combining these results with Assumption 5 and applying Lemma B.1, we obtain a bound on the overall Lipschitz constant.
(d) $-\frac{\partial h}{\partial x}\Bigr{\rvert}_{x=x}(f(x,g_{\bar{\theta},\bar{b}}(x,w)))-\mu_{h}h(x)$
$=-\frac{\partial h}{\partial x}\Bigr{\rvert}_{x=x}(f(x,g_{\bar{\theta},\bar{b}}(x,w)))-\mu_{h}h(x)+\mu_{h}h(x_{q})-\mu_{h}h(x_{q})+\frac{\partial h}{\partial x}\Bigr{\rvert}_{x=x_{q}}(f(x_{q},g_{\bar{\theta},\bar{b}}(x_{q},w_{q})))$
$\quad-\frac{\partial h}{\partial x}\Bigr{\rvert}_{x=x_{q}}(f(x_{q},g_{\bar{\theta},\bar{b}}(x_{q},w_{q})))$
$\leq\eta_{S}^{*}+(\mathcal{M}_{f}\mathcal{L}_{dh}+\mathcal{M}_{h}(\mathcal{L}_{x}+\sqrt{2}\mathcal{L}_{u}\mathcal{L}_{C}))\varepsilon+\mu_{h}\mathcal{L}_{h}\varepsilon$
$\leq\eta_{S}^{*}+\big{(}\mathcal{M}_{f}\mathcal{L}_{dh}+\mathcal{M}_{h}(\mathcal{L}_{x}+\sqrt{2}\mathcal{L}_{u}\mathcal{L}_{C})+\mu_{h}\mathcal{L}_{h}\big{)}\varepsilon\leq\eta_{S}^{*}+\mathcal{L}\varepsilon\leq 0.$
Therefore, if the condition (10) is satisfied, the neural $\delta$-ISS-CLF will ensure the system is incrementally input-to-state stable under the action of the controller. This completes the proof.
Appendix C Proof of Theorem 4.2

### v2

@@ 5.4 Jet Engine
We consider a nonlinear Moore-Grietzer Jet Engine Model in no-stall mode [jet_engine], whose dynamics is governed by the following set of equations.
@@
where $\mathsf{x}_{1}=\mu-1,\mathsf{x}_{2}=\zeta-\rho-2$ with $\mu,\zeta,\rho$ denote the mass flow, the pressure rise and a constant respectively. We consider the state space of the system to be ${\mathbf{X}}=[-0.25,0.25]\times[-0.25,0.25]$. Moreover, we consider the input set to be bounded within ${\mathbf{W}}=[-0.25,0.25]$. Also, we consider the model to be unknown. However, we estimate the Lipschitz constants $\mathcal{L}_{x}=1.213,\mathcal{L}_{u}=1$.
@@ 5.5 Rotating Spacecraft Model
We consider another example of a rotating rigid spacecraft model [khalil2002nonlinear], whose discrete-time dynamics is governed by the following set of equations.
@@
where $\mathsf{x}=[\mathsf{x}_{1},\mathsf{x}_{2},\mathsf{x}_{3}]^{\top}$ denotes angular velocities $\omega_{1},\omega_{2},\omega_{3}$ along the principal axes respectively, $\mathsf{u}=[\mathsf{u}_{1},\mathsf{u}_{2},\mathsf{u}_{3}]^{\top}$ represents the torque input, and with $J_{1}=200,J_{2}=200,J_{3}=100$ denote the principal moments of inertia. We consider the state space of the system to be ${\mathbf{X}}=[-0.25,0.25]\times[-0.25,0.25]\times[-0.25,0.25]$. Moreover, we consider the input set to be bounded within ${\mathbf{W}}=[-10,10]$. Note that the internal input has 3 dimensions, while the external input has a single dimension, that is, ${\mathbf{U}}\subset{\mathbb{R}}^{3}$, but $\mathbf{W}$$\subset{\mathbb{R}}$. In addition, we consider the model to be unknown. However, we estimate the Lipschitz constants $\mathcal{L}_{x}=0.160,\mathcal{L}_{u}=0.015$.
@@
One can see from Figure 5(a), 5(b), and 5(c), under the different input conditions, the trajectories corresponding to various states starting from different initial conditions maintain the same distance after some time instances under the influence of the controller.
@@ 5.6 Discussion and Comparisons
The proposed approach to synthesizing controllers, ensuring incremental stability for closed-loop continuous-time systems, provides several benefits. The data-driven neural controller synthesis utilizes the data from the state-space of the system. Existing approaches for designing controllers that ensure incremental input-to-state stability, such as those in [zamani2013backstepping, jagtap2017backstepping], require full knowledge of the system dynamics and assume a specific structure, typically a strict-feedback form. A recent work [sundarsingh2024backstepping] proposed estimating the strict feedback nature using the Gaussian process and designing a backstepping-like control to make the system incrementally stable. Compared to these methods, our approach can handle any general nonaffine nonlinear system, as seen in the Case Study 5.1 of a general nonlinear system. Although recent advancement [zaker2024controller] uses data collected from system trajectories to design a controller for an input-affine nonlinear system with polynomial dynamics, our method extends to arbitrary nonlinear dynamics, as seen in the first case study, where neither input-affine structures nor polynomial forms are present.
The final case study demonstrates that the system possesses three control signals, but can be effectively controlled using a reduced number of external inputs. Thereby, the proposed methodology is able to make the closed-loop system controllable with a reduced number of inputs, simplifying the control implementation. This will be particularly beneficial in improving the scalability of abstraction-based controller synthesis approaches [sadek2023compositional, symbolic1, jagtap2017quest]. As discussed in [zamani2017towards], the computational complexity of constructing an input-based abstraction of an incrementally stable system depends on the input dimension of the system. As we are able to make the system incrementally stable with a reduced number of input dimensions, the complexity of the abstraction is significantly reduced. This facilitates the efficient synthesis of additional formally verified controllers to meet other specifications for the closed-loop system.
@@ 6 Conclusion and Future Work
[…]
@@
The proof is inspired by the proof of Theorem 2.6 of [zamani2011lyapunov]. Let $\mathsf{x}_{x,\mathsf{w}}(t)$ denotes the state of the closed-loop system at time $t$ under input signal $\mathsf{w}$ from the initial condition $x=\mathsf{x}_{x,\mathsf{w}}(0)$. Note that, since the system is forward invariant under the action of the controller $g_{\bar{\theta},\bar{b}}$, for all $t\geq 0$ we have $\mathsf{x}_{x,\mathsf{w}}(t)\in{\mathbf{X}}$. Consider the closed-loop system admits a $\delta$-ISS-CLF under a controller. Now, by using property (i) of Definition 2.5, we get
@@
for any $t\in{\mathbb{R}}_{0}^{+}$. Now using condition (ii) of Definition 2.5 and the comparison lemma [khalil2002nonlinear, Lemma 3.4], we get
@@
$)\displaystyle\leq\alpha_{1}^{-1}\bigg(e^{-\kappa t}V(x,\hat{x})+\frac{1-e^{-\kappa t}}{\kappa}\sigma(\lVert\mathsf{w}-\hat{\mathsf{w}}\rVert$$\bigg)$
$)\displaystyle\leq\alpha_{1}^{-1}\bigg(e^{-\kappa t}V(x,\hat{x})+\frac{1}{\kappa}\sigma(\lVert\mathsf{w}-\hat{\mathsf{w}}\rVert$$\bigg)$
@@
[…]
$),\displaystyle|\mathsf{x}_{x,\mathsf{w}}(t)-\mathsf{x}_{\hat{x},\hat{\mathsf{w}}}(t)|\leq\xi(e^{-\kappa t}\alpha_{2}(|x-\hat{x}|))+\xi(\frac{1}{\kappa}\sigma(\lVert\mathsf{w}-\hat{\mathsf{w}}\rVert$)),
@@
$|\mathsf{x}_{x,\mathsf{w}}(t)-\mathsf{x}_{\hat{x},\hat{\mathsf{w}}}(t)|$ $).\displaystyle\leq\alpha_{1}^{-1}(2e^{-\kappa t}\alpha_{2}(|x-\hat{x}|))+\alpha_{1}^{-1}(\frac{2}{\kappa}\sigma(\lVert\mathsf{w}-\hat{\mathsf{w}}\rVert$)).
@@
[…]
$),\displaystyle\gamma(\lVert\mathsf{w}-\hat{\mathsf{w}}\rVert):=\alpha_{1}^{-1}(\frac{2}{\kappa}\sigma(\lVert\mathsf{w}-\hat{\mathsf{w}}\rVert$)), (23)
@@
Appendix B Proof of Lemma 2.8
Now, we assume that there exists a controller $g:{\mathbf{X}}\rightarrow{\mathbf{U}}$ such that condition (5) is satisfied. Then, one can infer at the boundary of $\mathbf{X}$}, the derivative of $h$ is always positive, enforcing the trajectories of the system under the controller $g$ to move towards inside the state space. Hence, the state space $\mathbf{X}$} becomes control forward invariant. This completes the proof.
Appendix C Proof of Lemma 3.3
[…]
$|\mathsf{L}_{f,g}V(x_{1},x_{2})-\mathsf{L}_{f,g}V(\hat{x}_{1},\hat{x}_{2})|\leq 2|\frac{\partial V_{\theta,b}}{\partial x_{1}}f(x_{1},g_{\theta,b}(x_{1},w_{1}))-\frac{\partial V_{\theta,b}}{\partial\hat{x}_{1}}f(\hat{x}_{1},g_{\theta,b}(\hat{x}_{1},\hat{w}_{1}))|\leq 2\big(\mathcal{M}_{f}\mathcal{L}_{dL}|x_{1}-\hat{x}_{1}|+\mathcal{M}_{L}|f(x_{1},g_{\theta,b}(x_{1},w_{1}))-f(\hat{x}_{1},g_{\theta,b}(\hat{x}_{1},\hat{w}_{1}))|\big)$.
Now, using Assumption 3 and 4, we have $|f(x_{1},g_{\theta,b}(x_{1},w_{1}))-f(\hat{x}_{1},g_{\theta,b}(\hat{x}_{1},\hat{w}_{1}))|\leq|f(x_{1},g_{\theta,b}(x_{1},w_{1}))-f(x_{1},g_{\theta,b}(x_{1},\hat{w}_{1}))+f(x_{1},g_{\theta,b}(x_{1},\hat{w}_{1}))-f(\hat{x}_{1},g_{\theta,b}(\hat{x}_{1},\hat{w}_{1}))|\leq(\mathcal{L}_{x}+\mathcal{L}_{u}\mathcal{L}_{c})|x_{1}-\hat{x}_{1}|+\mathcal{L}_{u}\mathcal{L}_{c}|w_{1}-\hat{w}_{1}|$.
[…]
Appendix D Proof of Theorem 3.4
With a slight abuse of notation, we denote $x_{\mathsf{w}}^{t}:=\mathsf{x}_{x,\mathsf{w}}(t)$ as the solution of the system (2) at time $t\geq 0$ under the input sequence $\mathsf{w}$ starting from the initial condition $x$. Therefore, we can write,
$V(x_{\mathsf{w}}^{\tau},\hat{x}_{\hat{\mathsf{w}}}^{\tau})=V(x,\hat{x})+\int_{0}^{\tau}\mathsf{L}_{f,g}V(x_{\mathsf{w}}^{t},\hat{x}_{\hat{\mathsf{w}}}^{t})dt.$ (24)
[…]
$\leq\frac{1}{\tau}\int_{0}^{\tau}|\mathsf{L}_{f,g}V(x_{\mathsf{w}}^{t},\hat{x}_{\hat{\mathsf{w}}}^{t})-\mathsf{L}_{f,g}V(x,\hat{x})|dt$
[…]
$\leq\frac{\mathscr{L}_{Vx}}{\tau}\int_{0}^{\tau}(|x_{\mathsf{w}}^{t}-x|+|\hat{x}_{\hat{\mathsf{w}}}^{t}-\hat{x}|)dt.$
Now we aim at finding the upper bound of $|x_{\mathsf{w}}^{t}-x|$. Under the continuity property of the solution of the system, we have $x_{\mathsf{w}}^{t}=x+\int_{0}^{t}f(x,g(x,\mathsf{w}(s)))ds$. Now, using Assumption 5, we can claim easily $|x_{\mathsf{w}}^{t}-x|\leq\int_{0}^{t}\mathcal{M}_{f}ds=\mathcal{M}_{f}t$. Similarly, we can also say, $|\hat{x}_{\hat{\mathsf{w}}}^{t}-\hat{x}|\leq\mathcal{M}_{f}t$. Therefore, replacing them back to the previous equation, we have
$|\widehat{\mathsf{L}}_{f,g}V(x,\hat{x})-\mathsf{L}_{f,g}V(x,\hat{x})|\leq\frac{\mathscr{L}_{Vx}}{\tau}\int_{0}^{\tau}\mathcal{M}_{f}tdt=\tau\mathscr{L}_{Vx}\mathcal{M}_{f},$
which completes the proof.
Appendix E Proof of Theorem 3.5
First, we prove that condition (9b) with $\hat{\eta}+\mathcal{L}\varepsilon\leq 0$ implies the satisfaction of condition (6b). Consider $x,\hat{x}\in{\mathbf{X}}$, then $V_{\theta,b}(x,\hat{x})-k_{1}|x-\hat{x}|^{\gamma_{1}}\leq V_{\theta,b}(x,\hat{x})-V_{\theta}(x_{q},x_{r})+V_{\theta}(x_{q},x_{r})-k_{1}|x-\hat{x}|^{\gamma_{1}}+k_{1}|x_{q}-x_{r}|^{\gamma_{1}}-k_{1}|x_{q}-x_{r}|^{\gamma_{1}}\leq\sqrt{2}\mathcal{L}_{L}\varepsilon+2\mathsf{L}_{1}\varepsilon+\hat{\eta}\leq\hat{\eta}+\mathcal{L}\varepsilon\leq 0$, showcasing the satisfaction of condition (6b). Similarly, one can follow these steps for the proof of satisfaction of other conditions, which implies that if we consider such $\hat{\eta}$ that satisfies $\hat{\eta}+\mathcal{L}\varepsilon\leq 0$, the lemma 3.1 has been satisfied, thereby establishing $V_{\theta,b}$ to be a valid $\delta$-ISS-CLF for the system (2) under the controller $g_{\bar{\theta},\bar{b}}$.
Appendix F Proof of Theorem 4.2